Huda opened a bank account with $30.00 and deposited $25.00 every month. After several months, she had a total of $180 in her account. For how many months did Huda save?

I know it's summation notation but the answer I am coming up with is not right and seems off. I thought it should be sigma and above it x and below it n=1 and then next to it 30(25)^n-1 so then the equation would be 30(1-25^x/1-25)=180 but I keep getting 1.546 and I know that is not right because it is not one of the options. Please help? thanks!

Answer by jim_thompson5910(35256)

Let x = number of months Huda saved

She starts with $30 and adds $25 each month. I'm assuming she deposits $25 each time. So for x months, she'll add on 25x dollars.

Total Saved Over x Months: 30 + 25x

This is set equal to 180 because "After several months, she had a total of $180 in her account"

---------------------------------------------------------------

30 + 25x = 180

25x = 180 - 30

25x = 150

x = 150/25

x = 6

She has saved for 6 months


Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim